# Phonon Induced Energy Relaxation in Quantum Critical Metals

### v1

@@ Abstract
Introduction.- In recent years, the problem of strange metals has drawn significant attention due to their unconventional transport behavior [1, 2, 3]. An experimental hallmark of these correlated metallic systems is a linear-in-temperature ($T$) electrical resistivity governed by a so-called Planckian relaxation rate of order $k_{B}T/\hbar$ down to low temperatures —in stark contrast to an umklapp-dominated $T^{2}$-scaling predicted by Fermi liquid theory, or a $T^{5}$-Bloch-Grüneissen regime due to scattering off acoustic phonons [4, 5]. The electrical transport timescale reflects the rate at which momentum is lost from the electronic system to external degrees of freedom, such as the lattice or impurities. Given the scale-invariant nature of the Planckian scattering rate [6] and experimental evidence for electronic quantum criticality across these materials [7, 8, 9, 10], a vast majority of theoretical efforts have focused on a scenario involving a Fermi surface coupled to the long-wavelength fluctuations of a bosonic collective mode of electronic origin [11, 12, 13], e.g. an electronic nematic [14]. The strong quantum critical fluctuations destroy the long-lived electronic quasiparticles near the Fermi surface, giving rise to non-Fermi liquid behavior. However, this does not automatically yield a $T-$linear momentum-relaxation rate, which has remained one of the central mysteries in the field. A number of theoretical works over the past few years have suggested additional microscopic ingredients that could potentially resolve this mystery [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29].
@@
Figure 1: (a) The bottleneck for energy relaxation from the electronic d.o.f. at local temperature $T_{e}$ to the phonons at local temperature $T_{p}$ is governed by $\Gamma_{ep}$, which is controlled by the processes illustrated in (b)-(d). The lowest-order Feynman diagrams for the phonon self-energy due to the couplings: (b) ${\mathcal{L}}_{ep}^{(1)}$ Eq. (4a), (c) ${\mathcal{L}}_{ep}^{(2)}$ Eq. (4b), and (d) ${\mathcal{L}}_{ep}^{(3)}$ Eq. (4c), respectively. Solid (dashed) lines denote renormalized electron (bosonic collective mode, $\varphi$) Green’s function.
@@
In this work, we focus on a related but even more poorly understood question: How does a strongly correlated metal relax its energy to the environment? While it is unlikely that phonons by themselves are responsible for the anomalous low-temperature momentum-relaxation in the strange-metal in the vicinity of quantum criticality [30], they are centrally important for energy relaxation (Fig. 1a). Phonons certainly act as a bath for the combined system consisting of electrons coupled to the bosonic collective mode, but have not received much theoretical attention in the context of non-Fermi liquids. The classic theoretical treatment [31] of energy relaxation rates, $\Gamma_{E}$, in a three-dimensional Fermi liquid coupled to acoustic phonons predicts a crossover from $\Gamma_{E}\propto T^{3}$ at low temperatures to $\Gamma_{E}\propto 1/T$ at high temperatures. In general, the momentum relaxation rate and $\Gamma_{E}$ are not automatically identical. How this behavior is modified in a non-Fermi liquid without long-lived electronic quasiparticles is presently unclear. We note that the theoretical framework developed below can be easily adapted to correlated insulating states that host Fermi surfaces of emergent neutral fermions (e.g. spinons) coupled to dynamical gauge-fields [32, 33], assuming that these degrees of freedom (d.o.f.) also relax their energy via coupling to acoustic phonons [34].
@@
Model of thermal relaxation.- We propose a minimal low-energy model consisting of three d.o.f.: electrons, a bosonic collective mode of electronic origin tuned to the vicinity of quantum criticality, and acoustic phonons (Fig. 1a). Our goal is to understand the temperature dependence of energy transfer from the combined electronic degrees of freedom to the phonons. The new ingredient relative to previous considerations is the presence of the bosonic collective mode, which we expect to couple to phonons via symmetry-allowed channels. The direct coupling between the bosonic collective mode and the phonons in the low-energy effective theory are associated with a distinct kinematic regime, compared to any usual local electron-phonon coupling. We investigate the effects of three distinct couplings: (i) a direct deformation potential type coupling between the electrons and acoustic phonons, and (ii) the two leading-order (i.e. a linear and quadratic) couplings between the bosonic collective mode and the acoustic phonons, each motivated by symmetry and microscopic considerations. In general, the electron-phonon coupling can lead to a modification of the electronic quantum critical point [35, 36, 37, 38], but these effects are restricted to a small fraction of the Brillouin-zone [38]. Since energy relaxation involves phonons of all momenta, we will ignore these modifications; we also assume that these interactions are weak compared to the dominant electronic interactions.
Before introducing the low-energy theory, we first describe a two-temperature model for thermal relaxation, which forms the basis for all the calculations. The electron-electron ($\Gamma_{ee}$) and the phonon-phonon ($\Gamma_{pp}$) energy relaxation rates are assumed to be much faster than the electron-phonon ($\Gamma_{ep}$) energy relaxation rate, respectively. The energy relaxation rate from the phonons to the environment ($\Gamma_{p-env}$) is assumed to be an even slower process, i.e. $\{\Gamma_{ee},\Gamma_{pp}\}\gg\Gamma_{ep}\gg\Gamma_{p-env}$. The energy equilibration proceeds as follows: First, the electrons and the phonons reach their individual equilibria over a short time scale set by $\Gamma_{ee}$ and $\Gamma_{pp}$, with two different temperatures $T_{e}$ and $T_{p}$, respectively.  [^1] For acoustic phonons at low temperatures, the assumption of $\Gamma_{pp}\gg\Gamma_{ep}$ may no longer be satisfied. This implies that phonons at different energies will equilibrate with the electrons at different speeds. If the electrons and phonons start with significantly different initial temperatures, the transient phonon distribution function will be far from equilibrium. However, if the electrons are only driven slightly out of equilibrium, we expect $T_{e}\approx T_{p}$. Then phonons at different energies still have approximately similar effective temperatures, and our theory treatment should still approximately remain valid. Second, the electrons and phonons equilibrate with each other via $\Gamma_{ep}$, reaching a temperature $T_{\text{f}}$. Finally, the system dissipates heat into the environment with relaxation rate $\Gamma_{p-env}$. Therefore, our theoretical treatment will assume that the electron sector and the phonon sector equilibrate rapidly to their respective (different) temperatures, and the bottleneck for energy-relaxation is limited by $\Gamma_{ep}$ (Fig. 1a). Assuming $T_{e}\approx T_{p}\approx T$, the relaxation process is described by,
$C_{e}\frac{\partial T_{e}}{\partial t}$ $=$ $-\kappa\left(T_{e}-T_{p}\right)\,,$ (1a)
$C_{p}\frac{\partial T_{p}}{\partial t}$ $=$ $\kappa\left(T_{e}-T_{p}\right)\,,$ (1b)
where $C_{e}$ and $C_{p}$ are heat capacities of the electrons and phonons, respectively, and $\kappa$ characterizes the energy flux between the electron and phonon sectors. We have ignored here the losses to the environment via $\Gamma_{p-env}$. The non-zero eigenvalue of Eq. (1) is denoted $-\Gamma_{E}$, which is precisely the energy relaxation rate of interest,
$\Gamma_{E}=\kappa\left(\frac{1}{C_{e}}+\frac{1}{C_{p}}\right)\,.$ (2)
@@
Note that the parametrically different dependence of $C_{e},~{}C_{p}$ on temperature inevitably accounts for some of the crossovers in $\Gamma_{E}$. To disentangle the thermodynamic $T-$dependent contributions to $C_{e},~{}C_{p}$ from the intrinsic $T-$dependence of the energy flux ($\kappa$) from electrons to phonons, in the remainder of this manuscript we focus primarily on $\kappa$ due to the distinct couplings described above. We will return to the full $T-$dependence of $\Gamma_{E}$ when we place our theoretical results in the context of the recent experiments [39].
@@
Here $c_{\bm{k}}^{\dagger}$ creates a fermion with momentum $\bm{k}$ and energy $\varepsilon_{\bm{k}}$, where $\sum_{\bm{k}}\equiv\int{\rm d}^{3}\bm{k}/(2\pi)^{3}$, and $\varphi_{\bm{{q}}}=\varphi_{-\bm{q}}^{\dagger}$ is the field operator for a quantum critical boson that carries no center of mass momentum (e.g. Ising-nematic order [14]). The boson velocity $v_{\varphi}$ is comparable to the Fermi velocity $v_{F}=\partial_{\bm{k}}\varepsilon_{\bm{k}}|_{k_{F}}$, and $r$ is used to tune to the quantum critical point (QCP) located at $r=r_{c}$. The Yukawa coupling of strength $g$ and form-factor $f_{\bm{k},\bm{q}}$ (e.g. $B_{1g}$ irreducible representation for the Ising-nematic) couples the electrons to the bosonic collective mode . The field $X_{i}$ denotes the $i$-th component of lattice displacement, and $c$ denotes the speed of sound. Here we have adopted the Debye model for phonons, i.e. the phonon momentum is cut off by ${\omega_{\rm D}}/c$, where ${\omega_{\rm D}}$ is the Debye frequency. For simplicity we have assumed that the different phonon polarizations have the same velocity. The ‘...’ in Eq. 3a represent a variety of generic perturbations, that can include disorder. Starting from Eq. (3a), a number of field-theoretic approaches yield a dynamical exponent $z_{\varphi}=3$ for the Landau-damped boson, and a non-Fermi liquid with $z_{c}=3/2$ in the clean theory [11, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 19, 72, 73, 74, 75, 17, 76, 20]. Inspired by a large body of experimental work [13], our goal is to focus on a scenario where in the absence of the coupling to phonons, the electrons are described by marginal Fermi liquid (MFL)-like correlations [77] in the quantum critical region. A variety of theoretical routes [43, 78], including disorder [71, 70, 19, 18], lead to $z_{\varphi}=2$, yielding an imaginary time MFL self-energy for the electrons, $\Sigma_{c}(i\omega)\sim-\omega\ln(\omega)$. While the exact microscopic mechanism leading to the ubiquitous evidence for MFL correlations is presently unclear, the following analysis of energy relaxation due to coupling to the acoustic phonons will be agnostic to the origin of MFL itself, as long as the electron-phonon interactions are weak.
We note that inspired in part by the recent experiments in cuprates, as well as other quasi two-dimensional correlated metals which display strange metallicity, here we will assume that while the electronic degrees of freedom are decoupled in the $z$-direction (i.e. the direct electronic hopping between layers is ignored), the phonons are entirely three-dimensional. We shall now include the three distinct channels for electron-phonon couplings, ${\mathcal{L}}_{ep}={\mathcal{L}}_{ep}^{(1)}+{\mathcal{L}}_{ep}^{(2)}+{\mathcal{L}}_{ep}^{(3)}$, where
${\mathcal{L}}_{ep}^{(1)}$ $=$ $\int{\rm d}\tau\sum_{\bm{k},\bm{q}}M^{i}_{\bm{k},\bm{q}}c^{\dagger}_{\bm{k+q/2}}c_{\bm{k-q/2}}X_{i,\bm{q}}\,,$ (4a)
${\mathcal{L}}_{ep}^{(2)}$ $=$ $\int{\rm d}\tau\sum_{\bm{q}}N_{\bm{q}}^{i}X_{i,-\bm{q}}\varphi_{\bm{q}}\,,$ (4b)
${\mathcal{L}}_{ep}^{(3)}$ $=$ $\frac{1}{2}\int{\rm d}\tau\sum_{\bm{k},\bm{{q}}}L^{i}_{\bm{k},\bm{q}}\varphi_{-\bm{k-q/2}}\varphi_{\bm{k+q/2}}X_{i,\bm{q}}\,.$ (4c)
Here, ${\mathcal{L}}_{ep}^{(1)}$ represents the standard electron-phonon coupling, where the matrix element $M^{i}_{\bm{k},\bm{q}}$ originates from coupling the lattice strain $(\partial_{i}X_{j}+\partial_{j}X_{i})$ to the electron density, and it encodes the correct linear combination in the symmetry allowed channel. ${\mathcal{L}}_{ep}^{(2)}$ is a direct linear coupling between the strain and the bosonic collective mode; such a coupling is allowed when the strain and $\varphi$ transform in the same fashion under crystalline symmetries, e.g. the nematic order can couple to the $B_{1g}$-strain modes [35]. Finally, ${\mathcal{L}}_{ep}^{(3)}$ represents the inelastic scattering process between the phonon and the bosonic collective mode [36]. According to Goldstone’s theorem, the matrix elements $M^{i}_{\bm{k},\bm{q}},N_{\bm{q}}^{i}$ and $L^{i}_{\bm{k},\bm{q}}$ are expected to be linear in the phonon momentum $\bm{q}$. For the coupling to electrons $M^{i}_{\bm{k},\bm{q}}$, the energy flux is neither sensitive to the specific form of $M^{i}_{\bm{k},\bm{q}}$ beyond the fact that $M^{i}_{\bm{k},\bm{q}}\propto\bm{q}$, nor to the quasi-two dimensional nature of the Fermi surface. However, for the critical boson $\varphi$, its quasi two-dimensional dispersion leads to important consequences. To be more concrete, the computation of the energy flux later relies on the matrix elements summed over polarizations, which we can be simplified based on Goldstone’s theorem $\{\sum_{i}|N^{i}_{\bm{q}}|,~{}\sum_{i}|L^{i}_{\bm{k},\bm{q}}|\}\sim A_{N,L}({\left|\bm{q_{\text{2D}}}\right|}^{2}+\lambda q_{z}^{2})$. Here, $A_{N}$ and $A_{L}$ are proportionality constants with the bare dimension of $[\text{energy}]^{2}$ and $[\text{energy}]^{0}$, respectively. The dependence on phonon momentum $\bm{q}$ can be decomposed into its 2D projection $\bm{q}_{\text{2D}}$ and the $z$-component $q_{z}$. The ${\left|\bm{q_{\text{2D}}}\right|}^{2}$ part arises from in-plane strain and the $q_{z}^{2}$ part arises from out-of-plane strain. We will assume that $\lambda\ll 1$ so that the $q_{z}^{2}$ term is a small contribution. We will comment later on the effects of larger $\lambda$.
As emphasized previously, we will assume that ${\mathcal{L}}_{ep}$ is weak compared to the self-interactions within the electron and the phonon sectors, allowing for a perturbative computation of the energy flux $\kappa$ between the two sectors. It is worth noting that the linear hybridization in ${\mathcal{L}}_{ep}^{(2)}$ can nominally be eliminated by an appropriate redefinition and change of basis of the $\varphi$-boson. However, we implicitly assume here that the phonons are responsible for heat dissipation to the environment. Eliminating Eq. (4b) via a basis transformation will then introduce a coupling between $\varphi$ and the environment, which we would like to avoid. Therefore, we explicitly include the weak linear hybridization in Eq. (4b) and keep track of how heat is dissipated away from the purely electronic components to the phonons, and subsequently to the environment.
A summary of our results is presented in Fig. 2, which are obtained using the Keldysh formalism [79]. The different contributions, $\kappa^{(\ell)}$, correspond to the energy flux due solely to the couplings, ${\mathcal{L}}_{ep}^{(\ell)}$ ($\ell=1,2,3$). All of these individual contributions add up together to yield the total $\kappa$, and $\Gamma_{E}$ is then determined via Eq. (2). In panel (a), we show the contribution due to the conventional electron-phonon coupling Eq. (4a), which is similar to Allen’s classic result [31] with a single crossover temperature scale around the Debye frequency ${\omega_{\rm D}}$. In panel (b), we show the additional structure in the $T$-dependent crossovers due to the boson-phonon couplings Eq. (4b) and (4c), which will be the subject of our discussion below.
@@
Figure 2: (a) Numerical plot of electronic contribution $\kappa^{(1)}$ to the energy flux. (b)Different regimes of the energy flux due to phonon-boson coupling (Eqs. (4b),(4c)) near the quantum critical point. $\omega_{D}$ is the Debye frequency of the phonons; $c$ is the speed of sound; $v_{\varphi}$ is the velocity of the $\varphi$ boson; $\gamma$ is the Landau damping coefficient (see Eq. (7)); $m_{0}$ is the renormalized boson mass which vanishes at the quantum critical point.
Energy-relaxation due to deformation potential coupling.- To make analytical progress in evaluating the energy flux, we have made some simplifying physical approximations, where we assume that (i) energy relaxation occurs homogeneously in space, (ii) the phonons remain reasonably well-defined quasiparticles, a point we revisit later, and in quasi-equilibrium at a temperature, $T_{p}$ [79]. The leading contribution due to ${\mathcal{L}}_{ep}^{(1)}$ in Eq. (4a) leads to the phonon self-energy in Fig. 1(b). When $T_{e}\approx T_{p}\approx T$, we obtain $\partial_{t}E=\kappa^{(1)}(T_{e}-T_{p})$, where the general expression for $\kappa^{(1)}$ is quite involved [79]. However, neglecting the electron self-energies and assuming $c\ll v_{F}$, we obtain
$\kappa_{\text{free}}^{(1)}=\frac{{\mathcal{N}}}{4v_{F}a_{z}}\int\frac{{\rm d}^{3}\bm{q}}{(2\pi)^{3}}\overline{|M({\left|\bm{q}\right|})|^{2}}\frac{c^{2}{\left|\bm{q}\right|}^{2}}{{\left|\bm{q_{\text{2D}}}\right|}}\frac{1}{T^{2}\sinh^{2}\frac{c{\left|\bm{q}\right|}}{2T}}\,,$ (5)
where $\bm{q}$ is the momentum transfer between the phonons and the particle-hole excitations, $a_{z}$ is the lattice constant in the $z-$direction and ${\mathcal{N}}$ is the electronic density of states. The $1/{\left|\bm{q_{\text{2D}}}\right|}$ factor arises from projecting onto the 2D Fermi surface, which however does not lead to any IR divergence. We have also replaced $\sum_{i}|M^{i}_{\bm{k}+\bm{q/2},\bm{q}}|^{2}$ by its average over the Fermi surface $\overline{|M(\bm{q})|^{2}}$. When $T\ll{\omega_{\rm D}}$ (Debye frequency), each factor of ${\left|\bm{q}\right|}$ and ${\left|\bm{q_{\text{2D}}}\right|}$ is replaced by $T$ (recall that $|M({\left|\bm{q}\right|})|^{2}\propto{\left|\bm{q}\right|}^{2}$), and we obtain $\kappa^{(1)}_{\text{free}}\propto T^{4}$. This result can also be alternatively derived using Fermi’s Golden rule, as follows:
$\begin{split}&\kappa^{(1)}\sim\overbrace{{\mathcal{N}}}^{\text{Fermion DOS}}\times\overbrace{T^{3}}^{\text{phonon density}}\times\overbrace{(\sqrt{T})^{2}}^{\text{Goldstone coupling}}\\ &\times\underbrace{T}_{\text{Energy transfer per particle}}\times\underbrace{\frac{1}{T}}_{\text{Expand around equilibrium}}\,.\end{split}$ (6)
Here a typical phonon has energy $T$ and the coupling enters as $\sqrt{T}$ because the displacement field $X$ is related to the creation and annihilation operators via $X_{\bm{q}}=(1/\sqrt{2c{\left|\bm{q}\right|}})(a_{\bm{q}}+a_{-\bm{q}}^{\dagger})$ [80]. On the other hand, when $T\gg{\omega_{\rm D}}$, Eq. (5) becomes $T$-independent, which is the classic expectation that in the equipartition regime of the phonons, it becomes difficult for them to absorb energy from the electrons. The numerical result of Eq. (5) is plotted in Fig. 2(a), which reproduces Allen’s result [31]; upon including the marginal Fermi liquid self-energy, the results remain nearly identical to the Fermi liquid results except at very low-temperatures [79].
@@
Energy-relaxation due to linear coupling between phonon and collective mode.- Let us now turn to the contribution due to ${\mathcal{L}}_{ep}^{(2)}$ in Eq. (4b), with a phonon self-energy shown in Fig. 1(c). As noted previously, the renormalized propagator for $\varphi$ in the MFL is overdamped with $z_{\varphi}=2$, with a spectral function,
@@
Here, $\gamma={\mathcal{N}}g^{2}/\Gamma$ is the Landau-damping coefficient with dimension of energy (${\mathcal{N}}$ is the fermion density of states, $\Gamma$ is the fermion elastic scattering rate), which we expect to be an electronic energy scale. The boson mass term, $m^{2}(T)$ also includes the thermal corrections due to a finite temperature, that are potentially important in the quantum critical regime. The QCP corresponds to $m^{2}(T=0)=0$, and the thermal fluctuations induce a $m^{2}(T)\sim\gamma T\ln(\gamma/T)$ [71, 81].

### v2

@@ Abstract
Introduction.- In recent years, the problem of strange metals has drawn significant attention due to their unconventional transport behavior [1, 2, 3]. An experimental hallmark of these correlated metallic systems is a linear-in-temperature ($T$) electrical resistivity governed by a so-called Planckian relaxation rate of order $k_{B}T/\hbar$ down to low temperatures —in stark contrast to an umklapp-dominated $T^{2}$-scaling predicted by Fermi liquid theory, or a $T^{5}$-Bloch-Grüneissen regime due to scattering off acoustic phonons [4, 5]. The electrical transport timescale reflects the rate at which momentum is lost from the electronic system to external degrees of freedom, such as the lattice or impurities. Given the scale-invariant nature of the Planckian scattering rate [6] and experimental evidence for electronic quantum criticality across these materials [7, 8, 9, 10], a vast majority of theoretical efforts have focused on a scenario involving a Fermi surface coupled to the long-wavelength fluctuations of a bosonic collective mode of electronic origin [11, 12, 13], e.g. an electronic nematic [14]. The strong quantum critical fluctuations destroy the long-lived electronic quasiparticles near the Fermi surface, giving rise to non-Fermi liquid behavior. Early important theoretical progress [15, 16, 17] already captured many experimental signatures of the strange metals in cuprates at a phenomenological level, especially the linear-in-$T$ single-particle scattering rate. However, this does not automatically yield a $T$-linear momentum-relaxation rate, which has remained one of the central mysteries in the field. A number of theoretical works over the past few years have suggested additional microscopic ingredients that could potentially resolve this mystery [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33].
@@
Figure 1: (a) The bottleneck for energy relaxation from the electronic d.o.f. at local temperature $T_{e}$ to the phonons at local temperature $T_{p}$ is governed by $\Gamma_{ep}$, which is controlled by the processes illustrated in (b)-(d). The lowest-order Feynman diagrams for the phonon self-energy due to the couplings: (b) ${\mathcal{L}}_{ep}^{(1)}$ Eq. (3), (c) ${\mathcal{L}}_{ep}^{(2)}$ Eq. (4), and (d) ${\mathcal{L}}_{ep}^{(3)}$ Eq. (5), respectively. Solid (dashed) lines denote renormalized electron (bosonic collective mode, $\varphi$) Green’s function.
@@
In this work, we focus on a related but even more poorly understood question: How does a strongly correlated metal relax its energy to the environment? Previous studies have focused a lot on how an electron becomes incoherent through inelastic scattering, but most of the transferred energy is assumed to remain in the electronic system. In this work, we query how this energy ultimately dissipates into the environment through various symmetry-allowed couplings to phonons. While it is unlikely that phonons by themselves are responsible for the anomalous low-temperature momentum-relaxation in the strange-metal in the vicinity of quantum criticality [34], they are centrally important for energy relaxation (Fig. 1a). Phonons certainly act as a bath for the combined system consisting of electrons coupled to the bosonic collective mode, but have not received much theoretical attention in the context of non-Fermi liquids. The classic theoretical treatment [35] of energy relaxation rates, $\Gamma_{E}$, in a three-dimensional Fermi liquid coupled to acoustic phonons predicts a crossover from $\Gamma_{E}\propto T^{3}$ at low temperatures to $\Gamma_{E}\propto 1/T$ at high temperatures. In general, the momentum relaxation rate and $\Gamma_{E}$ are not automatically identical. How this behavior is modified in a non-Fermi liquid without long-lived electronic quasiparticles is presently unclear. While our results are broadly applicable to a variety of correlated electron systems in the vicinity of quantum phase transitions, it is partly inspired by the quantum critical scenario in the cuprates, which remains an active area of investigation [36, 37, 38, 39, 40, 41, 42, 43]; see also alternative points of view [44, 45, 46, 47, 48]. We note that the theoretical framework developed below can be easily adapted to correlated insulating states that host Fermi surfaces of emergent neutral fermions (e.g. spinons) coupled to dynamical gauge-fields [49, 50], assuming that these degrees of freedom (d.o.f.) also relax their energy via coupling to acoustic phonons [51].
@@
Model of thermal relaxation.- We propose a minimal low-energy model consisting of three d.o.f.: electrons, a bosonic collective mode of electronic origin tuned to the vicinity of quantum criticality, and acoustic phonons (Fig. 1a). Our goal is to understand the temperature dependence of energy transfer from the combined electronic degrees of freedom to the phonons. The new ingredient relative to previous considerations is the presence of the bosonic collective mode, which we expect to couple to phonons via symmetry-allowed channels. The direct coupling between the bosonic collective mode and the phonons in the low-energy effective theory are associated with a distinct kinematic regime, compared to any usual local electron-phonon coupling. We investigate the effects of three distinct couplings: (i) a conventional electron–phonon deformation potential, and (ii) linear and (iii) quadratic couplings between the bosonic mode and phonons, each motivated by symmetry and microscopic considerations. We thus treat all couplings as weak perturbations compared to dominant electronic interactions.
To model the thermal relaxation process, we use the standard two-temperature model (illustrated in Fig. 1) which ignores the coupling between the phonons and the environment $\Gamma_{\text{p-env}}$, and assumes that the electrons and the phonons scattering rapidly within themselves and separately equilibrate first (with temperatures $T_{e}$ and $T_{p}$ respectively) and the relaxation bottleneck is set by the electron-phonon coupling. This is summarized by the hierarchy of scattering rates $\Gamma_{ee},\Gamma_{pp}\gg\Gamma_{ep}\gg\Gamma_{\text{p-env}}$ (see [52] for further discussion). The energy-flux between the electron and the phonon sector is given by $C_{p}{\rm d}T_{p}/{\rm d}t=-C_{e}{\rm d}T_{e}/{\rm d}t=\kappa(T_{e}-T_{p})$, and the energy relaxation rate is given by
$\Gamma_{E}=\kappa\left(\frac{1}{C_{e}}+\frac{1}{C_{p}}\right)\,$, (1)
@@
where $C_{e(p)}$ is the specific heat of the eletrons (phonons). Note that the parametrically different dependence of $C_{e},~C_{p}$ on temperature inevitably accounts for some of the crossovers in $\Gamma_{E}$. To disentangle the thermodynamic $T$-dependent contributions to $C_{e},~C_{p}$ from the intrinsic $T$-dependence of the energy flux ($\kappa$) from electrons to phonons, in the remainder of this manuscript we focus primarily on $\kappa$ due to the distinct couplings described above. In this work, we compute $\kappa$ using the Keldysh kinetic equation [52, 53, 54], and an alternative approach based on the Kubo formula is developed in [55]. We will return to the full $T$-dependence of $\Gamma_{E}$ when we place our results in the context of the recent experiments [56].
@@
Here, $c_{{\bm{k}}}^{\dagger}$ creates a fermion with momentum $\bm{k}$} and dispersion $\varepsilon_{{\bm{k}}}$, and $\varphi_{{\bm{q}}}=\varphi_{-{\bm{q}}}^{\dagger}$ denotes a critical bosonic mode (e.g., Ising-nematic order [14]) near a quantum critical point (QCP) tuned by $r$. Microscopically, the $\varphi$ field arises from Hubbard-Stratonovich decoupling of the electron interaction. The boson velocity $v_{\varphi}$ is comparable to the Fermi velocity $v_{F}$, and it couples to fermions via a Yukawa coupling $g$ and form factor $f_{{\bm{k}},{\bm{q}}}$ (e.g., $B_{1g}$ irreducible representation for the Ising-nematic). $X_{i}$ represents the $i$-th component of lattice displacement, with sound speed $c$, and we adopt a Debye phonon model with cutoff $\omega_{D}/c$. For simplicity, all phonon modes are taken to have the same velocity. The $\dots$ in Eq. (2a) includes generic perturbations such as disorder.
In the clean limit, the theory yields a Landau-damped boson with dynamical exponent $z_{\varphi}=3$ and a non-Fermi liquid with $z_{c}=3/2$ [11, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 22, 89, 90, 91, 92, 20, 93, 23]. Motivated by extensive experimental evidence [13], we focus on a regime where electrons display marginal Fermi liquid (MFL) behavior [15], with self-energy $\Sigma_{c}(i\omega)\sim-\omega\ln\omega$. This can arise from various mechanisms including disorder [88, 22, 21], which reduce the boson exponent to $z_{\varphi}=2$. Importantly, our analysis of energy relaxation via phonons is agnostic to the microscopic origin of MFL, assuming only that electron–phonon couplings are weak.
Motivated by experiments in cuprates and other quasi-two-dimensional strange metals, we assume that the electronic degrees of freedom are effectively two-dimensional, with negligible interlayer hopping, while phonons remain fully three-dimensional. We consider three symmetry-allowed electron-phonon couplings, ${\mathcal{L}}_{ep}={\mathcal{L}}_{ep}^{(1)}+{\mathcal{L}}_{ep}^{(2)}+{\mathcal{L}}_{ep}^{(3)}$, where:
${\mathcal{L}}_{ep}^{(1)}$ $=$ $\int{\rm d}\tau\sum_{{\bm{k}},{\bm{q}}}M^{i}_{{\bm{k}},{\bm{q}}}c^{\dagger}_{{\bm{k+q/2}}}c_{{\bm{k-q/2}}}X_{i,{\bm{q}}}\,$, (3)
${\mathcal{L}}_{ep}^{(2)}$ $=$ $\int{\rm d}\tau\sum_{{\bm{q}}}N_{{\bm{q}}}^{i}X_{i,-{\bm{q}}}\varphi_{{\bm{q}}}\,$, (4)
${\mathcal{L}}_{ep}^{(3)}$ $=$ $\frac{1}{2}\int{\rm d}\tau\sum_{{\bm{k}},{\bm{{q}}}}L^{i}_{{\bm{k}},{\bm{q}}}\varphi_{-{\bm{k-q/2}}}\varphi_{{\bm{k+q/2}}}X_{i,{\bm{q}}}\,.$ (5)
Here, ${\mathcal{L}}_{ep}^{(1)}$ is the standard deformation-potential coupling, where $M^{i}_{{\bm{k}},{\bm{q}}}$ encodes strain coupling to the electron density. ${\mathcal{L}}_{ep}^{(2)}$ describes a linear coupling between strain and the bosonic collective mode $\varphi$, allowed by symmetry (e.g., nematic modes and $B_{1g}$ strain [94]). ${\mathcal{L}}_{ep}^{(3)}$ captures inelastic scattering between phonons and pairs of $\varphi$ bosons [95]. By Goldstone’s theorem, the matrix elements $M^{i}_{{\bm{k}},{\bm{q}}},N^{i}_{{\bm{q}}},L^{i}_{{\bm{k}},{\bm{q}}}$ scale linearly with the phonon momentum $\bm{q}$}. The theory introduced above can be formally controlled using an extension of the large-$N$ Yukawa-SYK model [3, 88, 22, 21, 96, 93, 23, 87, 33], which we review in the End Matter.
For the electron coupling $M^{i}_{{\bm{k}},{\bm{q}}}$, only the $\bm{q}$}-scaling is relevant; the details of the Fermi surface geometry are largely unimportant. In contrast, the quasi-2D nature of $\varphi$ strongly affects the boson-phonon couplings. Summing over polarizations, we model the matrix elements as $\sum_{i}|N^{i}_{{\bm{q}}}|^{2},\sum_{i}|L^{i}_{{\bm{k}},{\bm{q}}}|^{2}\sim A_{N,L}({\left|{\bm{q_{\text{2D}}}}\right|}^{2}+\lambda q_{z}^{2})$, where $A_{N}$ and $A_{L}$ have dimensions $\text{energy}]^{2}$ and $\text{energy}]^{0}$, respectively. We assume $\lambda\ll 1$ to reflect weak coupling to out-of-plane strain, and comment later on its effects.
Throughout, we assume the couplings in ${\mathcal{L}}_{ep}$ are weak compared to electronic or phononic self-interactions, justifying a perturbative treatment of the energy flux $\kappa$. While ${\mathcal{L}}_{ep}^{(2)}$ could in principle be eliminated via a change of basis for $\varphi$, we retain it explicitly, since we treat phonons as the primary channel for energy dissipation into the environment. Removing it would spuriously couple $\varphi$ directly to the bath.
The contribution $\kappa^{(1)}$ from the deformation-potential electron-phonon coupling reproduces Allen’s well-known result, i.e. a crossover from $T^{4}$ to constant near the Debye frequency $\omega_{D}$ [35]. We review this in the End Matter and also demonstrate that the MFL self-energy does not change this picture. Fig. 2 summarizes the contributions $\kappa^{(2,3)}$ from the new boson-phonon couplings proposed here, which shows a richer $T$-dependent structure that we analyze below.
@@
Figure 2: Different regimes of the energy flux due to phonon-boson coupling (Eqs. (4),(5)) near the QCP. $\omega_{D}$ is the Debye frequency of the phonons; $c$ is the speed of sound; $v_{\varphi}$ is the velocity of the $\varphi$ boson; $\gamma$ is the Landau damping coefficient (see Eq. (6)); $m_{0}$ is the renormalized boson mass which vanishes at the QCP.
@@
Energy-relaxation due to linear coupling between phonon and collective mode.- Let us now turn to the contribution due to ${\mathcal{L}}_{ep}^{(2)}$ in Eq. (4), with a phonon self-energy shown in Fig. 1(c). As noted previously, the renormalized propagator for $\varphi$ in the MFL is overdamped with $z_{\varphi}=2$, with a spectral function,
@@
Here, $\gamma={\mathcal{N}}g^{2}/\Gamma$ is the Landau-damping coefficient with dimension of energy ($\mathcal{N}$}: fermion density of states, $\Gamma$: fermion elastic scattering rate), expected to be an electronic energy scale. The boson mass term, $m^{2}(T)$ includes the corrections due to a finite temperature, that are potentially important in the quantum critical regime. The QCP corresponds to $m(T=0)=0$, and the thermal fluctuations induce a $m^{2}(T)\sim\gamma T\ln(\gamma/T)$ [88, 97].